OpenStudy (anonymous):

integrate 1/((t^3)(sqrt((t^2)-4)) from 2sqrt2 to 4

OpenStudy (anonymous):

by the bounds i bet $12 it is an inverse trig answer, even though i can't read it

OpenStudy (anonymous):

I already know it is, i need to substitute in using sec^2(theta)-1=tan^2(theta), but it keeps telling me my answer is wrong

OpenStudy (anonymous):

\[\int\limits_{2\sqrt{2}}^{4}\frac{ 1 }{t ^{3}\sqrt{t ^{2}-4}}\] is the equation

OpenStudy (tkhunny):

It's not an equation. It is an integral expression. I'm pretty tempted to rewrite the argument: \(\dfrac{t}{t^{4}\sqrt{t^{2}-4}}\) and try some astute substitutions.